Order: 2/3 strength sustacal 300 ml p.o. q.i.d

Supply: sustacal 5 and 10 fl oz cans
Answer #1

Order: 2/3 strength sustacal 300ml p.o q.i.d

1 oz= 30ml

5 oz = 5X30 = 150ml

10 oz = 10X30 = 300ml

300 ml q.i.d (Four times, 300X4) = 1200ml

2/3 strength =1200 X2/3 = 800

= 1200-800 = 400 ml

Therefore 800ml of medication and 400ml of water/ diluent to be added.

If 10 oz can is used, 2 full cans + 200ml from another container can be used.

If 5 oz can is used, 5 full cans + 50 ml from another container can be used.

If both 10oz and 5 oz cans are used, 2 10oz cans, 1 5oz can and 50ml from 5oz can could be used.